# Dynamical black hole entropy beyond general relativity from the Einstein frame

### v2

@@ Abstract
Recently Hollands, Wald and Zhang proposed a new formula for the entropy of a dynamical black hole for an arbitrary theory of gravity obtained from a diffeomorphism covariant Lagrangian via the Noether charge methodHollands:2024vbe . We present an alternative, pedagogical derivation of the dynamical black hole entropy for $f(R)$ gravity as well as canonical scalar-tensor theory by means of conformal transformations. First, in general relativity we generalize Visser and Yan’s pedagogical proof of the non-stationary physical process first law to black holes that may not be in vacuum, and give a pedagogical derivation of the second-order behavior of the dynamical black hole entropy for vacuum perturbations by considering the second-order variation of the Raychaudhuri equation. Second, we apply the derivation for general relativity to theories in the Einstein frames, and then recast the conclusions in their original frames. We show that the dynamical black hole entropy formulae of these theories satisfy both the non-stationary physical process first law and the non-stationary comparison first law via the Einstein frame. We further study the second-order behavior of the dynamical black hole entropy for vacuum perturbations, and observe that the second law is obeyed at second order in those theories. Using the Einstein frame, we also determine the relationship between the dynamical black hole entropy and the Wald entropy of the generalized apparent horizon in the original frame.
Keywords:
Black Holes, Classical Theories of Gravity
1 Introduction
The discovery of the laws of black hole thermodynamics is one of the great achievements of fundamental physicsBardeen:1973gs ; Bekenstein:1973ur ; Hawking:1975vcx . These laws combine gravity, quantum theory and thermodynamics within one stunning framework. Moreover, they provide some of the deepest insights on the fundamental nature of the quantum theory of gravity. For general relativity (GR), the black hole entropy $S$ is given by the Bekenstein-Hawking formulaBekenstein:1973ur ; Hawking:1975vcx
@@ 1 Introduction
where $A$ is the area of the event horizon and $G$ is Newton’s constant. While for an arbitrary theory of gravity obtained from a diffeomorphism covariant Lagrangian, the entropy of a stationary black hole at the bifurcation surface $\mathcal{B}$ is given by the Wald entropy, which is defined as the Noether charge and for the $f(\text{Riemann})$ theories of gravity it is written asWald:1993nt ; Iyer:1994ys
@@
where $v$ is the future-directed affine parameter of the null generator of the future horizon, $u$ denotes the affine null distance away from the horizon and is also future-directed. Furthermore, Iyer and Wald proposed a formula $S_{\text{Iyer-Wald}}$ for dynamical black hole entropy on an arbitrary cross-section $\mathcal{C}$ of the horizon by expanding the Wald entropy in fields and their derivatives, and keeping the terms that depend only on boost-invariant fieldsIyer:1994ys . However, the second law of black hole mechanics does not seem to hold for $S_{\text{Iyer-Wald}}$. And that proposed dynamical entropy formula is not field redefinition invariant. Moreover, the Noether charge method used to derive the black hole entropy is subjected to a number of ambiguities for non-stationary black holes, identified by Jacobson, Kang and Myers (JKM)Jacobson:1993vj . Wall derived a second law for higher curvature gravity in a perturbative context, and resolved some ambiguities in Wald’s Noether charge methodWall:2015raa . For $f(\text{Riemann})$ gravity the Wall entropy is given by
$S_{\text{Wall}}=-8\pi\int_{\mathcal{C}(v)}\mathrm{d}A\left(\frac{\partial L}{\partial R_{uvuv}}-4\frac{\partial^{2}L}{\partial R_{uiuj}\partial R_{vkvl}}K_{ij(u)}K_{kl(v)}\right).$ (3)
@@
Although the second law for $S_{\text{BH}}$ and the linearized second law for $S_{\text{Wall}}$ hold for non-stationary perturbations, this is not the case for the first law. The first law often does not hold for non-stationary perturbations of a stationary black hole, and if it does, the black hole entropy cannot be evaluated at an arbitrary cross-section of the event horizon of the perturbed non-stationary black holeVisser:2024pwz . Recently, Hollands, Wald and Zhang proposed a strategy to overcome these two limitationsHollands:2024vbe . They derived the first law by applying the Noether charge method to non-stationary perturbations of a stationary black hole background, and introduced a new program to the definition of dynamical black hole entropy valid to leading order, on the basis of the validity of a local, “physical process version” of the first law of black hole mechanics. Visser and Yan generalized and improved their work in a number of ways, and gave a more pedagogical proof of the physical process first law for black holes in general relativityVisser:2024pwz .
Next let us review the key results of Hollands:2024vbe for the dynamical black hole entropy and the non-stationary first law. For a non-stationary dynamical black hole, the entropy $S$ is no longer equal to the Bekenstein-Hawking entropy for GR or to the Wall entropy for higher curvature gravity. The formula for dynamical black holes differs from the usual Noether charge formula by a non-trivial dynamical correction term. For general relativity, the dynamical black hole entropy was defined as
@@
where $v$ is the future-directed affine null parameter along the future horizon, and it is equal to 0 at the bifurcation surface $\mathcal{B}$. Notably, the derivative term is invariant under the scaling transformation of the affine parameter $v\to a(x^{i})v$, where $x^{i}$ are codimension-2 spatial coordinates on the horizon. It was also shown that, to leading order in perturbation theory, the formula of dynamical black hole entropy for general relativity is equal to the Bekenstein-Hawking entropy of the apparent horizon. Furthermore, the dynamical black hole entropy was generalized to higher curvature gravity asHollands:2024vbe
@@
where $M$ and $J$ are the mass and angular momentum of the black hole, respectively. $\Omega_{\mathcal{H}}$ is the angular velocity of the horizon. And $\kappa$ denotes the surface gravity of the black hole. For vacuum perturbations of a stationary black hole, the dynamical entropy is “time independent” to first order, i.e., $\delta S_{\text{dyn}}[\mathcal{C}(v)]=\delta S_{\text{dyn}}[\mathcal{B}]$, where $\mathcal{C}(v)$ is an arbitrary cross-section. In order to study the non-trivial time variation of black hole entropy, we may unseal an external stress-energy, $\delta T_{ab}$, in the first-order perturbation. For perturbations sourced by external matter fields, the non-stationary physical process first law reads [^1] Note that in the traditional treatments of the physical process first law, the black hole starts and ends in a stationary stateGao:2001ut ; Poisson:2009pwt .
$\frac{\kappa}{2\pi}\Delta\delta S_{\text{dyn}}=\int_{v_{1}}^{v_{2}}\mathrm{d}v\int_{\mathcal{C}(v)}\mathrm{d}A\ \kappa v\delta T_{vv}=\Delta\delta M-\Omega_{\mathcal{H}}\Delta\delta J,$ (7)
@@
In this paper we present an alternative, pedagogical proof of both the non-stationary physical process first law and the non-stationary comparison first law for the $f(R)$ gravity as well as the canonical scalar-tensor theory via conformal transformations. To begin with, we generalize Visser and Yan’s pedagogical derivation of the physical process first law for GR to black holes that may not be in vacuum. We also present a pedagogical deviation of the second-order behavior of $S_{\text{dyn}}$ for vacuum perturbations in GR by studying the second variation of the Raychaudhuri equation. Since conformal transformations preserve the causal structures of spacetimes, and the conformal factors that put theories into the Einstein frames are independent of the Killing time $\tau$ in the stationary spacetimes, the spacetimes given by conformal transformations are stationary black hole solutions to the Einstein equation. We are capable of applying the proof of the first law for GR to theories in the Einstein frames, and then recast them in their original frames. As the perturbations are non-stationary, the gauge conditions for perturbations that fix the affine parameter $v^{E}$ in the Einstein frame will not necessarily fix the affine parameter $v$ in the original frame. However, those gauge conditions do not restrict the way we perturb the stationary black hole background. Instead, they tell us how to identify spacetime points in the two slightly different spacetimes. We show that the switching from $v^{E}-$identification to $v-$identification leads to second-order corrections, and the first laws are invariant to first order under the re-identification of spacetime points. For vacuum perturbations we further study the second-order behavior of the dynamical black hole entropy, and observe that the second law holds at second order both in the Einstein frames and in the original frames. Using the Einstein frame we also determine the relationship between the dynamical black hole entropy of the cross-section and the Wald entropy of the generalized apparent horizon in the original frame, which was not given before.
@@
The rest of the paper is organized as follows. In section 2 we introduce the geometric setup and impose gauge conditions on perturbations. In Section 3 we generalize Visser and Yan’s pedagogical proof of the physical process first law for GR to non-vacuum solutions, and study the second-order behavior of the dynamical black hole entropy for vacuum perturbations. In Sections 4 and 5 we present the proof of both the physical process first law and the comparison first law for those theories, study the second-order behavior of the dynamical black hole entropy for vacuum perturbations, and determine the relationship between the dynamical black hole entropy and the Wald entropy of the generalized apparent horizon in the original frame. Section 6 is devoted to the summary of the paper and some discussions. In Appendix A we calculate the modified canonical energy flux for $f(R)$ gravity via the covariant phase space formalism, and give a non-trivial check of the second-order behavior of $S_{\text{dyn}}$ for vacuum perturbations.
@@
We will mainly follow the notation and conventions of Wald:1984rg . Throughout this paper, we set $c=\hbar=k_{B}=1$ while keep Newton’s constant $G_{D}$ explicit.
2 Stationary black hole background and gauge conditions for perturbations
@@
In this section we introduce the geometry of the stationary black hole background and impose gauge conditions on non-stationary perturbations used to study the first law. Our geometric setup mainly follows that of Visser:2024pwz .
$\mathcal{H}^{+}$$\mathcal{H}^{-}$$k^{a}$$l^{a}$$\mathcal{B}$$\mathcal{C}(v)$$\xi^{a}$
Figure 1: The stationary black hole background
@@
Consider a stationary $D-$dimensional stationary black hole background geometry $(M,g_{ab})$. We assume that the black hole spacetime is asymptotic flat and electrically neutral. The event horizon of this black hole coincides with the bifurcate Killing Horizon $\mathcal{H}$. We label the future horizon by $\mathcal{H}^{+}$, the past horizon by $\mathcal{H}^{-}$, and the bifurcation surface by $\mathcal{B}$. The Killing field normal to the Killing horizon is denoted by $\xi^{a}$, and it is a Killing symmetry of the metric $g_{ab}$ as well as the matter fields $\phi$. i.e.,
@@
We would like to perturb the stationary black hole background $g_{ab}\to g_{ab}+\delta g_{ab}$ and the matter fields $\phi\to\phi+\delta\phi$ to study the first law for non-stationary variations, where $\delta g_{ab},\ \delta\phi\sim\mathcal{O}(\epsilon)$ are of first order in the perturbation. Since the stationary black hole spacetime and the perturbed spacetime are different, there exists certain gauge freedoms about how we should establish the coordinate systems in the perturbed geometry and how to identify spacetime points in the two slightly different spacetimes. In order to simplify the deviation of the first law, we impose the following gauge conditions on perturbationsVisser:2024pwz :
@@
    and $k^{a}$ remains null normal to $\mathcal{H}^{+}$ and $l^{a}$ remains null everywhere under the perturbation, which together with the condition $\delta(k^{a}l_{a})=0$ yields the following conditions
@@
    The Killing vector field $\xi^{a}$ remains null and tangent to the geodesic generators of the perturbed black hole. Combining with $\xi^{a}\delta g_{ab}\overset{\mathcal{H}}{=}0$ means that $\delta\xi^{a}$ is proportional to $k^{a}$ on $\mathcal{H}^{+}$ and to $l^{a}$ on $\mathcal{H}^{-}$.
@@
3 Dynamical black hole entropy in general relativity
In this section, we present the formula for dynamical black hole entropy based on the non-stationary physical process first law in GR. This derivation is completed by means of integrating the linearized Raychaudhuri equation on the horizon between two arbitrary cross-sections. A similar derivation was given in Visser:2024pwz for black hole solutions to the vacuum Einstein equation. While our deviation is less restrictive since we allow the metric to couple to matter fields via the Einstein equation in the stationary background. For vacuum perturbations we also provide a pedagogical derivation of the second-order behavior of the dynamical black hole entropy by considering the second-order variation of the Raychaudhuri equation. We also note that a similar deduction for vacuum perturbations was established in Rignon-Bret:2023fjq before us.
@@ 3 Dynamical black hole entropy in general relativity
Where $G_{D}$ is the $D-$dimensional Newton’s constant. As a corollary of the Raychaudhuri equation combined with the Einstein equation, $T_{ab}k^{a}k^{b}$ is equal to zero on the horizon in the stationary geometry. Otherwise the last term on the right hand side of (14) would not be zeroPoisson:2009pwt . Physically this condition can be interpreted as matter cannot be flowing across the event horizon. We multiply the Raychaudhuri equation on both sides by $\kappa v$, integrate it over the future horizon between cross-sections $\mathcal{C}(v_{1})$ and $\mathcal{C}(v_{2})$, and then vary this equation. Since $\frac{\mathrm{d}\theta}{\mathrm{d}v}$ and $T_{ab}k^{a}k^{b}$ vanish in the unperturbed geometry, $\delta$ acts only on $\theta$ on the left-hand side and only on $R_{ab}$ on the right-hand side, and the right-hand side is left with $8\pi G_{D}\kappa v\delta T_{ab}k^{a}k^{b}$ due to our gauge condition $\delta g_{ab}k^{a}k^{b}=0$. Recalling that $\xi^{a}=\kappa vk^{a}$ on the future horizon, to first order the result is
@@
where $\mathrm{d}A=\mathrm{d}^{D-2}x\sqrt{\gamma(x,v)}$ is the area element of the cross-section $\mathcal{C}(v)$. Let $x=(x^{1},\cdots x^{D-2})$ be a coordinate system on $\mathcal{C}(v_{1})$, we can extend this coordinate system to an arbitrary cross-section by requiring that all points on each null geodesic share the same values of $(x^{1},\cdots x^{D-2})$. To exchange the order of integrals in (16), define the function $\mathcal{A}(x,v)$ on $\mathcal{H}^{+}$ asFlanagan:1999jp
$\mathcal{A}(x,v)=\exp\left[\int_{v_{1}}^{v}\mathrm{d}\tilde{v}\ \theta(x,\tilde{v})\right].$ (17)
The integral on the left hand side of (16) can then be split into
$\int_{v_{1}}^{v_{2}}\mathrm{d}v\int_{\mathcal{C}(v)}\mathrm{d}A\ v\frac{\mathrm{d}\delta\theta}{\mathrm{d}v}=\int_{\mathcal{C}(v_{1})}\mathrm{d}A\int_{v_{1}}^{v_{2}}\mathrm{d}v\ \mathcal{A}(x,v)v\frac{\mathrm{d}\delta\theta}{\mathrm{d}v},$ (18)
@@
$\begin{split}\delta\int_{\mathcal{C}(v_{1})}\mathrm{d}A\int_{v_{1}}^{v_{2}}\mathrm{d}v\ \theta&=\delta\int_{\mathcal{C}(v_{1})}\mathrm{d}A\int_{v_{1}}^{v_{2}}\mathrm{d}v\ \mathcal{A}(x,v)\partial_{v}\ln\sqrt{\gamma(x,v)}\\ &=\delta\int_{\mathcal{C}(v_{1})}\mathrm{d}^{D-2}x\sqrt{\gamma(x,v_{1})}\int_{v_{1}}^{v_{2}}\mathrm{d}v\ \frac{\sqrt{\gamma(x,v)}}{\sqrt{\gamma(x,v_{1})}}\frac{\partial_{v}\sqrt{\gamma(x,v)}}{\sqrt{\gamma(x,v)}}\\ &=\delta\int_{\mathcal{C}(v_{1})}\mathrm{d}^{D-2}x\left[\sqrt{\gamma(x,v_{2})}-\sqrt{\gamma(x,v_{1})}\right]\\ &=\Delta\delta A.\end{split}$ (20)
@@
And $\Delta$ denotes the difference between $\mathcal{C}(v_{1})$ and $\mathcal{C}(v_{2})$. By the argument similar to (16), if we vary $\Delta E$, $\delta$ acts only on $T_{ab}$. It was shown by Hawking that if a black hole is stationary, then it must be either static or axisymmetricHawking:1971vc . If the horizon Killing field is normalized as $\xi^{a}=\left(\partial_{t}\right)^{a}+\Omega_{\mathcal{H}}\left(\partial_{\theta}\right)^{a}$, where $\left(\partial_{t}\right)^{a}$ represents time translations at infinity, $\left(\partial_{\theta}\right)^{a}$ represents the rotational Killing vector and $\Omega_{\mathcal{H}}$ is the angular velocity of the black hole, then the mass and angular momentum transferred across the horizon arePoisson:2009pwt

### v4

@@ Abstract
Recently Hollands, Wald and Zhang proposed a new formula for the entropy of a dynamical black hole for an arbitrary theory of gravity obtained from a diffeomorphism covariant Lagrangian via the Noether charge method[1]. We present an alternative, pedagogical derivation of the dynamical black hole entropy for $f(R)$ gravity as well as canonical scalar-tensor theory by means of conformal transformations. First, in general relativity we generalize Visser and Yan’s pedagogical proof of the non-stationary physical process first law to black holes that may not be in vacuum, and give a pedagogical derivation of the second-order behavior of the dynamical black hole entropy for vacuum perturbations by considering the second-order variation of the Raychaudhuri equation. Second, we apply the derivation for general relativity to theories in the Einstein frames, and then recast the conclusions in their original frames. We show that the dynamical black hole entropy formulas of these theories satisfy both the non-stationary physical process first law and the non-stationary comparison first law via the Einstein frame. We further study the second-order behavior of the dynamical black hole entropy for vacuum perturbations, and observe that the second law is obeyed at second order in those theories. Using the Einstein frame, we also determine the relationship between the dynamical black hole entropy and the Wald entropy of the generalized apparent horizon in the original frame.
I Introduction
The discovery of the laws of black hole thermodynamics is one of the great achievements of fundamental physics[2, 3, 4]. These laws combine gravity, quantum theory and thermodynamics within one stunning framework. Moreover, they provide some of the deepest insights on the fundamental nature of the quantum theory of gravity. For general relativity (GR), the black hole entropy $S$ is given by the Bekenstein-Hawking formula[3, 4]
@@ I Introduction
where $A$ is the area of the event horizon and $G$ is Newton’s constant. While for an arbitrary theory of gravity obtained from a diffeomorphism covariant Lagrangian, the entropy of a stationary black hole at the bifurcation surface $\mathcal{B}$ is given by the Wald entropy, which is defined as the Noether charge and for the $f(\text{Riemann})$ theories of gravity it is written as[5, 6]
@@
where $v$ is the future-directed affine parameter of the null generator of the future horizon, $u$ denotes the affine null distance away from the horizon and is also future-directed. Furthermore, Iyer and Wald proposed a formula $S_{\text{Iyer-Wald}}$ for dynamical black hole entropy on an arbitrary cross-section $\mathcal{C}$ of the horizon by expanding the Wald entropy in fields and their derivatives, and keeping the terms that depend only on boost-invariant fields[6]. However, the second law of black hole mechanics does not seem to hold for $S_{\text{Iyer-Wald}}$. And that proposed dynamical entropy formula is not field-redefinition invariant. Moreover, the Noether charge method used to derive the black hole entropy is subjected to a number of ambiguities for non-stationary black holes, identified by Jacobson, Kang and Myers (JKM)[7]. Wall derived a second law for higher curvature gravity in a perturbative context, and resolved some ambiguities in Wald’s Noether charge method[8]. For $f(\text{Riemann})$ gravity the Wall entropy is given by
$S_{\text{Wall}}=-8\pi\int_{\mathcal{C}(v)}\mathrm{d}A\left(\frac{\partial L}{\partial R_{uvuv}}\right.\left.-4\frac{\partial^{2}L}{\partial R_{uiuj}\partial R_{vkvl}}K_{ij(u)}K_{kl(v)}\right).$ (3)
@@
Although the second law for $S_{\text{BH}}$ and the linearized second law for $S_{\text{Wall}}$ hold for non-stationary perturbations, this is not the case for the first law. The first law often does not hold for non-stationary perturbations of a stationary black hole, and if it does, the black hole entropy cannot be evaluated at an arbitrary cross-section of the event horizon of the perturbed non-stationary black hole[9]. Recently, Hollands, Wald and Zhang proposed a strategy to overcome these two limitations[1]. They derived the first law by applying the Noether charge method to non-stationary perturbations of a stationary black hole background, and introduced a new program to the definition of dynamical black hole entropy valid to leading order, on the basis of the validity of a local, “physical process version” of the first law of black hole mechanics. Visser and Yan generalized and improved their work in a number of ways, and gave a more pedagogical proof of the physical process first law for black holes in general relativity[9].
Next let us review the key results of [1] for the dynamical black hole entropy and the non-stationary first law. For a non-stationary dynamical black hole, the entropy $S$ is no longer equal to the Bekenstein-Hawking entropy for GR or to the Wall entropy for higher curvature gravity. The formula for dynamical black holes differs from the usual Noether charge formula by a nontrivial dynamical correction term. For general relativity, the dynamical black hole entropy was defined as
@@
where $v$ is the future-directed affine null parameter along the future horizon, and it is equal to 0 at the bifurcation surface $\mathcal{B}$. Notably, the derivative term is invariant under the scaling transformation of the affine parameter $v\to a(x^{i})v$, where $x^{i}$ are codimension-2 spatial coordinates on the horizon. It was also shown that, to leading order in perturbation theory, the formula of dynamical black hole entropy for general relativity is equal to the Bekenstein-Hawking entropy of the apparent horizon. Furthermore, the dynamical black hole entropy was generalized to higher curvature gravity as[1]
@@
where $M$ and $J$ are the mass and angular momentum of the black hole, respectively. $\Omega_{\mathcal{H}}$ is the angular velocity of the horizon. And $\kappa$ denotes the surface gravity of the black hole. For vacuum perturbations of a stationary black hole, the dynamical entropy is “time independent” to first order, i.e., $\delta S_{\text{dyn}}[\mathcal{C}(v)]=\delta S_{\text{dyn}}[\mathcal{B}]$, where $\mathcal{C}(v)$ is an arbitrary cross-section. In order to study the nontrivial time variation of black hole entropy, we may unseal an external stress-energy, $\delta T_{ab}$, in the first-order perturbation. For perturbations sourced by external matter fields, the non-stationary physical process first law reads [^1] Note that in the traditional treatments of the physical process first law, the black hole starts and ends in a stationary state[10, 11].
$\begin{split}\frac{\kappa}{2\pi}\Delta\delta S_{\text{dyn}}&=\int_{v_{1}}^{v_{2}}\mathrm{d}v\int_{\mathcal{C}(v)}\mathrm{d}A\ \kappa v\delta T_{vv}=\Delta\delta M-\Omega_{\mathcal{H}}\Delta\delta J,\end{split}$ (7)
@@
In this paper we present an alternative, pedagogical proof of both the non-stationary physical process first law and the non-stationary comparison first law for the $f(R)$ gravity as well as the canonical scalar-tensor theory via conformal transformations. To begin with, we generalize Visser and Yan’s pedagogical derivation of the physical process first law for GR to black holes that may not be in vacuum. We also present a pedagogical derivation of the second-order behavior of $S_{\text{dyn}}$ for vacuum perturbations in GR by studying the second variation of the Raychaudhuri equation. Since conformal transformations preserve the causal structures of spacetimes, and the conformal factors that put theories into the Einstein frames are independent of the Killing time $\tau$ in the stationary spacetimes, the spacetimes given by conformal transformations are stationary black hole solutions to the Einstein equation. We are capable of applying the proof of the first law for GR to theories in the Einstein frames, and then recast them in their original frames. As the perturbations are non-stationary, the gauge conditions for perturbations that fix the affine parameter $v^{E}$ in the Einstein frame will not necessarily fix the affine parameter $v$ in the original frame. However, those gauge conditions do not restrict the way we perturb the stationary black hole background. Instead, they tell us how to identify spacetime points in the two slightly different spacetimes. We show that the switching from $v^{E}-$identification to $v-$identification leads to second-order corrections, and the first laws are invariant to first order under the reidentification of spacetime points. For vacuum perturbations we further study the second-order behavior of the dynamical black hole entropy, and observe that the second law holds at second order both in the Einstein frames and in the original frames. Using the Einstein frame we also determine the relationship between the dynamical black hole entropy of the cross-section and the Wald entropy of the generalized apparent horizon in the original frame, which was not given before.
@@
The rest of the paper is organized as follows. In Sec. II we introduce the geometric setup and impose gauge conditions on perturbations. In Sec. III we generalize Visser and Yan’s pedagogical proof of the physical process first law to non-vacuum solutions, and study the second-order behavior of the dynamical black hole entropy for vacuum perturbations in GR. In Secs. IV and V we present the proof of both the physical process first law and the comparison first law for $f(R)$ gravity and canonical scalar-tensor theory, study the second-order behavior of the entropy for vacuum perturbations in those theories, and determine the relationship between the dynamical black hole entropy and the Wald entropy of the generalized apparent horizon in the original frame. Sec. VI is devoted to the summary of the paper and some discussions. In Appendix A we calculate the modified canonical energy flux for $f(R)$ gravity via the covariant phase space formalism, and give a nontrivial check of the second-order behavior of $S_{\text{dyn}}$ for vacuum perturbations.
@@
We will mainly follow the notation and conventions of [12]. Throughout this paper, we set $c=\hbar=k_{B}=1$ while keep Newton’s constant $G_{D}$ explicit.
II Stationary black hole background and gauge conditions for perturbations
@@
In this section we introduce the geometry of the stationary black hole background and impose gauge conditions on non-stationary perturbations used to study the first law. Our geometric setup mainly follows that of [9].
Figure 1: The stationary black hole background.
@@
Consider a $D-$dimensional stationary black hole background geometry $(M,g_{ab})$, as shown in FIG. 1. We assume that the black hole spacetime is asymptotically flat and electrically neutral. The event horizon of this black hole coincides with the bifurcate Killing Horizon $\mathcal{H}$. We label the future horizon by $\mathcal{H}^{+}$, the past horizon by $\mathcal{H}^{-}$, and the bifurcation surface by $\mathcal{B}$. The Killing field normal to the Killing horizon is denoted by $\xi^{a}$, and it is a Killing symmetry of the metric $g_{ab}$ as well as the matter fields $\phi$. i.e.,
@@
We would like to perturb the stationary black hole background $g_{ab}\to g_{ab}+\delta g_{ab}$ and the matter fields $\phi\to\phi+\delta\phi$ to study the first law for non-stationary variations, where $\delta g_{ab},\ \delta\phi\sim\mathcal{O}(\epsilon)$ are of first order in the perturbation. Since the stationary black hole spacetime and the perturbed spacetime are different, there exists certain gauge freedoms about how we should establish the coordinate systems in the perturbed geometry and how to identify spacetime points in the two slightly different spacetimes. In order to simplify the derivation of the first law, we impose the following gauge conditions on perturbations[9]:
@@
    and $k^{a}$ remains null normal to $\mathcal{H}^{+}$ and $l^{a}$ remains null everywhere under the perturbation. Combined with the requirement $\delta(k^{a}l_{a})=0$ this implies the following conditions
@@
    The Killing vector field $\xi^{a}$ remains null and tangent to the geodesic generators of the perturbed black hole. Combined with $\xi^{a}\delta g_{ab}\overset{\mathcal{H}}{=}0$ this means that $\delta\xi^{a}$ is proportional to $k^{a}$ on $\mathcal{H}^{+}$ and to $l^{a}$ on $\mathcal{H}^{-}$.
@@
III Dynamical black hole entropy in general relativity
In this section, we present the formula for dynamical black hole entropy based on the non-stationary physical process first law in GR. This derivation is completed by means of integrating the linearized Raychaudhuri equation on the horizon between two arbitrary cross-sections. A similar derivation was given in [9] for black hole solutions to the vacuum Einstein equation. While our derivation is less restrictive since we allow the metric to couple to matter fields via the Einstein equation in the stationary background. For vacuum perturbations we also provide a pedagogical derivation of the second-order behavior of the dynamical black hole entropy by considering the second-order variation of the Raychaudhuri equation. We also note that a similar deduction for vacuum perturbations was established in [13] before us.
@@ III Dynamical black hole entropy in general relativity
Where $G_{D}$ is the $D-$dimensional Newton’s constant. As a corollary of the Raychaudhuri equation combined with the Einstein equation, $T_{ab}k^{a}k^{b}$ is equal to zero on the horizon in the stationary geometry. Otherwise the last term on the right-hand side of (14) would not be zero[11]. Physically this condition can be interpreted as matter cannot be flowing across the event horizon. We multiply the Raychaudhuri equation on both sides by $\kappa v$, integrate it over the future horizon between cross-sections $\mathcal{C}(v_{1})$ and $\mathcal{C}(v_{2})$, and then vary this equation. Since $\frac{\mathrm{d}\theta}{\mathrm{d}v}$ and $T_{ab}k^{a}k^{b}$ vanish in the unperturbed geometry, $\delta$ acts only on $\theta$ on the left-hand side and only on $R_{ab}$ on the right-hand side, and the right-hand side is left with $8\pi G_{D}\kappa v\delta T_{ab}k^{a}k^{b}$ due to our gauge condition $\delta g_{ab}k^{a}k^{b}=0$. Recalling that $\xi^{a}=\kappa vk^{a}$ on the future horizon, to first order the result is
@@
where $\mathrm{d}A=\mathrm{d}^{D-2}x\sqrt{\gamma(x,v)}$ is the area element of the cross-section $\mathcal{C}(v)$. Let $x=(x^{1},\cdots x^{D-2})$ be a coordinate system on $\mathcal{C}(v_{1})$, we can extend this coordinate system to an arbitrary cross-section by requiring that all points on each null geodesic share the same values of $(x^{1},\cdots x^{D-2})$. To exchange the order of integrals in (16), define the function $\mathcal{A}(x,v)$ on $\mathcal{H}^{+}$ as[14]
$\mathcal{A}(x,v)=\exp\left[\int_{v_{1}}^{v}\mathrm{d}\tilde{v}\ \theta(x,\tilde{v})\right]=\frac{\sqrt{\gamma(x,v)}}{\sqrt{\gamma(x,v_{1})}}.$ (17)
The integral on the left-hand side of (16) can then be split into
$\begin{split}\int_{v_{1}}^{v_{2}}\mathrm{d}v\int_{\mathcal{C}(v)}\mathrm{d}A\ v\frac{\mathrm{d}\delta\theta}{\mathrm{d}v}&=\int_{v_{1}}^{v_{2}}\mathrm{d}v\int_{\mathcal{C}}\mathrm{d}^{D-2}x\sqrt{\gamma(x,v)}v\frac{\mathrm{d}\delta\theta}{\mathrm{d}v}\\ &=\int_{v_{1}}^{v_{2}}\mathrm{d}v\int_{\mathcal{C}}\mathrm{d}^{D-2}x\sqrt{\gamma(x,v_{1})}\frac{\sqrt{\gamma(x,v)}}{\sqrt{\gamma(x,v_{1})}}v\frac{\mathrm{d}\delta\theta}{\mathrm{d}v}\\ &=\int_{\mathcal{C}}\mathrm{d}^{D-2}x\sqrt{\gamma(x,v_{1})}\int_{v_{1}}^{v_{2}}\mathrm{d}v\frac{\sqrt{\gamma(x,v)}}{\sqrt{\gamma(x,v_{1})}}v\frac{\mathrm{d}\delta\theta}{\mathrm{d}v}\\ &=\int_{\mathcal{C}(v_{1})}\mathrm{d}A\int_{v_{1}}^{v_{2}}\mathrm{d}v\ \mathcal{A}(x,v)v\frac{\mathrm{d}\delta\theta}{\mathrm{d}v},\end{split}$ (18)
@@
$\begin{split}&\delta\int_{\mathcal{C}(v_{1})}\mathrm{d}A\int_{v_{1}}^{v_{2}}\mathrm{d}v\ \theta\\ =&\delta\int_{\mathcal{C}(v_{1})}\mathrm{d}A\int_{v_{1}}^{v_{2}}\mathrm{d}v\ \mathcal{A}(x,v)\partial_{v}\ln\sqrt{\gamma(x,v)}\\ =&\delta\int_{\mathcal{C}(v_{1})}\mathrm{d}^{D-2}x\sqrt{\gamma(x,v_{1})}\int_{v_{1}}^{v_{2}}\mathrm{d}v\ \frac{\sqrt{\gamma(x,v)}}{\sqrt{\gamma(x,v_{1})}}\frac{\partial_{v}\sqrt{\gamma(x,v)}}{\sqrt{\gamma(x,v)}}\\ =&\delta\int_{\mathcal{C}(v_{1})}\mathrm{d}^{D-2}x\left[\sqrt{\gamma(x,v_{2})}-\sqrt{\gamma(x,v_{1})}\right]\\ =&\Delta\delta A.\end{split}$ (20)
@@
And $\Delta$ denotes the difference between $\mathcal{C}(v_{1})$ and $\mathcal{C}(v_{2})$. By the argument similar to (16), if we vary $\Delta E$, $\delta$ acts only on $T_{ab}$. It was shown by Hawking that if a black hole is stationary, then it must be either static or axisymmetric[15]. If the horizon Killing field is normalized as $\xi^{a}=\left(\partial_{t}\right)^{a}+\Omega_{\mathcal{H}}\left(\partial_{\theta}\right)^{a}$, where $\left(\partial_{t}\right)^{a}$ represents time translations at infinity, $\left(\partial_{\theta}\right)^{a}$ represents the rotational Killing vector and $\Omega_{\mathcal{H}}$ is the angular velocity of the black hole, then the mass and angular momentum transferred across the horizon are[11]